Compton Corporation is a wholesale distributor of educational CD-ROMs. The company's records indicate the following:

This Year Last Year
Units Sold 250,000 200,000
Sales $ 1,250,000 $ 1,000,000
Cost of goods sold 875,000 700,000
Gross margin 375,000 300,000
Selling and administrative expenses 222,000 210,000
Net operating income $ 153,000 $ 90,000

What is the company's contribution margin for this year?
Answer #1

Calculate contribution margin this year :

Variable cost per unit selling and administrative expense = 12000/50000 = 0.24 per unit

Contribution margin = Sales-Variable cost

= 1250000-(3.74*250000)

Contribution margin = 315000